# TerraCodec: Compressing Earth Observations

### v1

@@ G.1 Quantitative results
Table 7: Test SSIM (↑), RMSE (↓), and MAE (↓) on AllClear across difficulty subsets (by average cloudiness). Following the benchmark, the metrics are computed across all bands, not per-band.
@@
Model SSIM RMSE MAE SSIM RMSE MAE SSIM RMSE MAE SSIM RMSE MAE
@@
CTGAN 0.765 0.084 0.072 0.794 0.068 0.056 0.822 0.056 0.044 0.840 0.052 0.041
DiffCR 0.716 0.075 0.063 0.739 0.071 0.061 0.758 0.068 0.059 0.744 0.068 0.060
PMAA 0.746 0.071 0.060 0.758 0.076 0.066 0.770 0.080 0.071 0.768 0.086 0.078
U-TILISE 0.546 0.254 0.226 0.598 0.211 0.185 0.693 0.153 0.134 0.807 0.097 0.083
UnCRtainTS 0.813 0.061 0.046 0.826 0.063 0.049 0.865 0.057 0.044 0.898 0.050 0.039
TerraCodec-$T$ 0.814 0.064 0.050 0.830 0.065 0.050 0.887 0.045 0.034 0.917 0.034 0.025
Table 7 reports SSIM, RMSE, and MAE results for all methods, complementing the PSNR results in the main paper (Table 1). Our zero-shot TEC–TT clearly outperform the heuristics: TEC–TT reaches PSNR $\approx 32.9$ dB and SSIM $\approx 0.917$, compared to LeastCloudy ($30.61$ dB / $0.863$) and Mosaicing ($29.82$ dB / $0.755$). Relative to the strongest prior zero-shot neural method on AllClear (UnCRtainTS, $29.01$ dB / $0.898$ / MAE =0.039=0.039 / RMSE =0.050=0.050), TEC–TT is substantially stronger ($32.86$ dB / $0.917$ / MAE =0.025=0.025 / RMSE =0.034=0.034).
@@
G.2 Qualitative results
Figure 19 presents additional TEC–TT inpainting results on the AllClear test set across varying degrees of cloudiness in the three past input frames. While three context images are shown, TEC–TT uses only the two least cloudy as input. Notably, even under heavily clouded conditions, TEC–TT leverages past context to produce relatively accurate predictions of the target frame.
Figure 19: Cloud inpainting examples with TEC–$T$ on the AllClear benchmark. Reflectance values are clipped to 0–2000, which causes clouds to appear saturated in the visualizations.

### v2

@@ 0.G.1 Quantitative results
Table 7: Test RMSE (↓) and MAE (↓) on AllClear across difficulty subsets (by average cloudiness). Metrics computed across all bands.
@@
Model RMSE MAE RMSE MAE RMSE MAE RMSE MAE
@@
CTGAN 0.084 0.072 0.068 0.056 0.056 0.044 0.052 0.041
DiffCR 0.075 0.063 0.071 0.061 0.068 0.059 0.068 0.060
PMAA 0.071 0.060 0.076 0.066 0.080 0.071 0.086 0.078
U-TILISE 0.254 0.226 0.211 0.185 0.153 0.134 0.097 0.083
UnCRtainTS 0.061 0.046 0.063 0.049 0.057 0.044 0.050 0.039
TerraCodec-$T$ 0.064 0.050 0.065 0.050 0.045 0.034 0.034 0.025
Table 7 reports RMSE, and MAE results for all methods, complementing the PSNR and SSIM results in the main paper zero-shot TEC-TT clearly outperform the heuristics: TEC-TT reaches PSNR $\approx 32.9$ dB and SSIM $\approx 0.917$, compared to LeastCloudy ($30.61$ dB / $0.863$) and Mosaicing ($29.82$ dB / $0.755$). Relative to the strongest prior zero-shot neural method on AllClear (UnCRtainTS, $29.01$ dB / $0.898$ / MAE =0.039=0.039 / RMSE =0.050=0.050), TEC-TT is substantially stronger ($32.86$ dB / $0.917$ / MAE =0.025=0.025 / RMSE =0.034=0.034).
@@
0.G.2 Qualitative results
Figure 19 presents additional TEC-TT inpainting results on the AllClear test set across varying degrees of cloudiness in the three past input frames. While three context images are shown, TEC-TT uses only the two least cloudy as input. Notably, even under heavily clouded conditions, TEC-TT leverages past context to produce relatively accurate predictions of the target frame.
Figure 19: Cloud inpainting examples with TEC-$T$ on the AllClear benchmark. Reflectance values are clipped to 0–2000, which causes clouds to appear saturated in the visualizations.
Use of LLMs